OpenStudy (anonymous):

What is the equation of the line that is perpendicular to -x + y = 7 and passes through (-1, -1)?

OpenStudy (anonymous):

a)-x + y = 2 b)-x – y = 2 c)x – y = 2 d)-x – y = 0 e)x + y = 0